OpenStudy (anonymous):

y=x^3 has a point of inflection at x=o. I know it is true but how i prove it with an statement?

OpenStudy (anonymous):

You test the point by, take the Second Derivato on it.. and then you can test it by using, a point before 0 and after.. For exempel test with the point -1 and 1

OpenStudy (anonymous):

you get a parabola

OpenStudy (anonymous):

and the parabola touches the origin..is this fine?

OpenStudy (anonymous):

@nabaz thankyou. it means i have to solve it.

OpenStudy (anonymous):

No problem, is it a parabola...???

OpenStudy (anonymous):

yeah , of course, u need proof?

OpenStudy (anonymous):

i thouht it was a Curve ..

OpenStudy (anonymous):

@thinker thankyou

OpenStudy (anonymous):

if so, u have the answer wid u...just try substituting x=-1,1,0,-2,2,...... and you'll get corresponding y values, with which u can get a curve...and that curve is a parabola

OpenStudy (anonymous):

now , is this fine?? @nabaz & @abhinn

OpenStudy (anonymous):

yeah, i last thing... u mean this is a parabola.. http://en.wikipedia.org/wiki/File:X_cubed_plot.svg

OpenStudy (anonymous):

Copie the text.. could get the hole adress..

OpenStudy (anonymous):

no no....not that one...:).try this http://en.wikipedia.org/wiki/File:Parabola.svg this is a parabola...both of go to this link n see

OpenStudy (anonymous):

*both of u

OpenStudy (anonymous):

its like a "U" shape

OpenStudy (anonymous):

Yes but this was a Y = x^3 and its not a parabole

OpenStudy (anonymous):

Yes, but x3 is like a S with is lying downwards.

OpenStudy (anonymous):

yeah...y=x^3 is a symmetric curve, but parabola looks like a "U" ...

OpenStudy (anonymous):

haha.. my bad, i thought u meant that x^3 was a parabola

OpenStudy (anonymous):

its ok:) nabaz

OpenStudy (anonymous):

i wud like to give u both a hint

OpenStudy (anonymous):

u der?

OpenStudy (anonymous):

hint for what?

OpenStudy (anonymous):

to know if an equation is parabola or what..only if u need it:)

OpenStudy (anonymous):

yes, i know i think. ever x^2 x^4 x^6 .. ( not odd nrs )

OpenStudy (anonymous):

exactly :)...good